Having problems solving for x and y on this one. Please help?
Solving by subsitution, either independent, dependent, or inconsistent?
y = -3x + 19 and y = 2x -1

Answer by tinbar(133)

1)y=-3x + 19
2)y=2x-1
so take he right side of the equation in number 2 and replace it for y in number 1
2x-1 = -3x+19
5x=20
x=4
now sub x=4 back into 1 or 2 to get y, both will give same answer
from 1)y=-3(4)+19 = -12+19 = 7
from 2)y=2(4)-1 = 8-1 = 7
therefore x=4, y =7